# All-in-One: Heterogeneous Interaction Modeling for Cold-Start Rating Prediction

### v2

@@ 17       Update model parameters by gradient descent;
UC NeuMF $0.3328_{\small(.0106)}$ $0.3887_{\small(.0110)}$ $0.2657_{\small(.0167)}$ $0.4176_{\small(.0112)}$ $0.3962_{\small(.0132)}$ $0.3388_{\small(.0203)}$ $0.5165_{\small(.0155)}$ $0.4178_{\small(.0127)}$ $0.4214_{\small(.0271)}$
Wide&Deep $0.2852_{\small(.0161)}$ $0.5408_{\small(.0197)}$ $0.2161_{\small(.0008)}$ $0.3759_{\small(.0231)}$ $0.5528_{\small(.0273)}$ $0.2517_{\small(.0196)}$ $0.4746_{\small(.0678)}$ $0.5666_{\small(.0416)}$ $0.3258_{\small(.0504)}$
DeepFM $0.2956_{\small(.0013)}$ $0.5154_{\small(.0066)}$ $0.1870_{\small(.0079)}$ $0.3511_{\small(.0125)}$ $0.5230_{\small(.0151)}$ $0.2120_{\small(.0070)}$ $0.4321_{\small(.0278)}$ $0.5224_{\small(.0015)}$ $0.2787_{\small(.0233)}$
AFN $0.2205_{\small(.0714)}$ $0.4970_{\small(.0354)}$ $0.1462_{\small(.0644)}$ $0.3195_{\small(.0414)}$ $0.5063_{\small(.0264)}$ $0.1856_{\small(.0452)}$ $0.4210_{\small(.0066)}$ $0.5238_{\small(.0080)}$ $0.2573_{\small(.0277)}$
MAMO $0.4016_{\small(.0115)}$ $0.2752_{\small(.0194)}$ $0.3062_{\small(.0123)}$ $0.5444_{\small(.0070)}$ $0.3244_{\small(.0184)}$ $0.4419_{\small(.0091)}$ $0.7119_{\small(.0046)}$ $0.3895_{\small(.0150)}$ $0.6417_{\small(.0065)}$
TaNP $0.4118_{\small(.0030)}$ $0.8504_{\small(.0145)}$ $0.3338_{\small(.0039)}$ $0.4598_{\small(.0054)}$ $0.8601_{\small(.0110)}$ $0.3626_{\small(.0131)}$ $0.5288_{\small(.0210)}$ $0.8712_{\small(.0072)}$ $0.4160_{\small(.0329)}$
MeLU $0.4651_{\small(.0416)}$ $0.5860_{\small(.0709)}$ $0.3534_{\small(.0440)}$ $0.5282_{\small(.0779)}$ $0.6079_{\small(.0737)}$ $0.4079_{\small(.0662)}$ $0.6701_{\small(.0442)}$ $0.6466_{\small(.0798)}$ $0.5558_{\small(.0492)}$
HIRE $\mathbf{0.5713_{\small(.0530)}}$ $\mathbf{0.8931_{\small(.0291)}}$ $\mathbf{0.5079_{\small(.0751)}}$ $\mathbf{0.6405_{\small(.0257)}}$ $\mathbf{0.9058_{\small(.0267)}}$ $\mathbf{0.5640_{\small(.0544)}}$ $\mathbf{0.7191_{\small(.0284)}}$ $\mathbf{0.9295_{\small(.0234)}}$ $\mathbf{0.6475_{\small(.0151)}}$
IC NeuMF $0.4070_{\small(.0066)}$ $0.3632_{\small(.0407)}$ $0.3282_{\small(.0414)}$ $0.5142_{\small(.0225)}$ $0.3813_{\small(.0422)}$ $0.4303_{\small(.0773)}$ $0.5930_{\small(.0868)}$ $0.4103_{\small(.0430)}$ $0.5729_{\small(.0200)}$
Wide&Deep $0.5007_{\small(.0786)}$ $0.8014_{\small(.0095)}$ $0.3814_{\small(.0935)}$ $0.5451_{\small(.0885)}$ $0.8059_{\small(.0096)}$ $0.3953_{\small(.0874)}$ $0.5870_{\small(.0914)}$ $0.8047_{\small(.0058)}$ $0.4393_{\small(.0928)}$
DeepFM $0.5246_{\small(.0813)}$ $0.8110_{\small(.0162)}$ $0.4092_{\small(.0772)}$ $0.5142_{\small(.0723)}$ $0.8156_{\small(.0181)}$ $0.4121_{\small(.0928)}$ $0.5515_{\small(.0769)}$ $0.8123_{\small(.0219)}$ $0.4437_{\small(.0995)}$
AFN $0.4915_{\small(.0710)}$ $0.8018_{\small(.0178)}$ $0.4040_{\small(.0908)}$ $0.5395_{\small(.0998)}$ $0.8083_{\small(.0187)}$ $0.3764_{\small(.0761)}$ $0.5679_{\small(.0660)}$ $0.8129_{\small(.0187)}$ $0.4399_{\small(.0891)}$
MAMO $0.4129_{\small(.0062)}$ $0.2810_{\small(.0049)}$ $0.3246_{\small(.0051)}$ $0.5462_{\small(.0041)}$ $0.3240_{\small(.0036)}$ $0.4498_{\small(.0071)}$ $0.7160_{\small(.0036)}$ $0.3942_{\small(.0086)}$ $0.6418_{\small(.0019)}$
TaNP $0.4116_{\small(.0315)}$ $0.8545_{\small(.0132)}$ $0.3125_{\small(.0152)}$ $0.4506_{\small(.0174)}$ $0.8618_{\small(.0129)}$ $0.3287_{\small(.0290)}$ $0.5194_{\small(.0435)}$ $0.8713_{\small(.0122)}$ $0.3848_{\small(.0504)}$
MeLU $0.4925_{\small(.0577)}$ $0.6159_{\small(.0289)}$ $0.3764_{\small(.0517)}$ $0.5204_{\small(.0820)}$ $0.6258_{\small(.0245)}$ $0.3915_{\small(.0639)}$ $0.6415_{\small(.0872)}$ $0.6565_{\small(.0362)}$ $0.5341_{\small(.0789)}$
HIRE $\mathbf{0.5837_{\small(.0338)}}$ $\mathbf{0.8925_{\small(.0056)}}$ $\mathbf{0.5174_{\small(.0327)}}$ $\mathbf{0.6418_{\small(.0244)}}$ $\mathbf{0.9057_{\small(.0057)}}$ $\mathbf{0.5649_{\small(.0206)}}$ $\mathbf{0.7220_{\small(.0471)}}$ $\mathbf{0.9262_{\small(.0065)}}$ $\mathbf{0.6519_{\small(.0497)}}$
U&I C NeuMF $0.3829_{\small(.0097)}$ $0.4221_{\small(.0329)}$ $0.2976_{\small(.0105)}$ $0.4342_{\small(.0120)}$ $0.4304_{\small(.0345)}$ $0.3274_{\small(.0229)}$ $0.4922_{\small(.0074)}$ $0.4364_{\small(.0205)}$ $0.3661_{\small(.0170)}$
Wide&Deep $0.4037_{\small(.0841)}$ $0.7387_{\small(.0776)}$ $0.3304_{\small(.0696)}$ $0.5023_{\small(.0682)}$ $0.7172_{\small(.0840)}$ $0.4058_{\small(.0692)}$ $0.5649_{\small(.0654)}$ $0.7204_{\small(.0790)}$ $0.3126_{\small(.0866)}$
DeepFM $0.3927_{\small(.0771)}$ $0.6848_{\small(.0731)}$ $0.3018_{\small(.0710)}$ $0.4750_{\small(.0294)}$ $0.6900_{\small(.0838)}$ $0.3649_{\small(.0429)}$ $0.5414_{\small(.0965)}$ $0.7266_{\small(.0503)}$ $0.3881_{\small(.0981)}$
AFN $0.3476_{\small(.0396)}$ $0.6344_{\small(.0479)}$ $0.2815_{\small(.0301)}$ $0.3505_{\small(.0193)}$ $0.6298_{\small(.0468)}$ $0.2651_{\small(.0085)}$ $0.3988_{\small(.0317)}$ $0.6458_{\small(.0479)}$ $0.2839_{\small(.0289)}$
MAMO $0.4100_{\small(.0087)}$ $0.3256_{\small(.0632)}$ $0.3026_{\small(.0376)}$ $0.5167_{\small(.0763)}$ $0.3699_{\small(.0534)}$ $0.4101_{\small(.0934)}$ $0.7178_{\small(.0154)}$ $0.4485_{\small(.0724)}$ $0.6380_{\small(.0391)}$
TaNP $0.5114_{\small(.0174)}$ $0.8812_{\small(.0650)}$ $0.4365_{\small(.0083)}$ $0.5153_{\small(.0214)}$ $0.8774_{\small(.0447)}$ $0.4286_{\small(.0456)}$ $0.5484_{\small(.0676)}$ $0.8780_{\small(.0250)}$ $0.4531_{\small(.0856)}$
MeLU $0.4335_{\small(.0436)}$ $0.5465_{\small(.0146)}$ $0.3349_{\small(.0293)}$ $0.4992_{\small(.0928)}$ $0.5716_{\small(.0080)}$ $0.3784_{\small(.0788)}$ $0.6373_{\small(.0918)}$ $0.5993_{\small(.0351)}$ $0.5350_{\small(.0999)}$
HIRE $\mathbf{0.6077_{\small(.0333)}}$ $\mathbf{0.9060_{\small(.0025)}}$ $\mathbf{0.5529_{\small(.0322)}}$ $\mathbf{0.6903_{\small(.0430)}}$ $\mathbf{0.9209_{\small(.0051)}}$ $\mathbf{0.6284_{\small(.0405)}}$ $\mathbf{0.7636_{\small(.0689)}}$ $\mathbf{0.9392_{\small(.0064)}}$ $\mathbf{0.7015_{\small(.0812)}}$

### v3

@@ 15       Update model parameters by gradient descent;
UC Popularity $37.68_{\small(1.07)}$ $83.77_{\small(0.39)}$ $26.33_{\small(1.55)}$ $42.33_{\small(1.74)}$ $84.14_{\small(0.19)}$ $27.51_{\small(1.69)}$ $48.01_{\small(0.14)}$ $86.10_{\small(0.21)}$ $31.29_{\small(0.95)}$
TransGNN $38.75_{\small(1.76)}$ $41.43_{\small(1.72)}$ $32.11_{\small(1.56)}$ $41.07_{\small(2.52)}$ $43.44_{\small(2.27)}$ $34.34_{\small(2.35)}$ $45.01_{\small(0.07)}$ $46.24_{\small(0.67)}$ $36.83_{\small(1.53)}$
NeuMF $33.28_{\small(1.06)}$ $38.87_{\small(1.10)}$ $26.57_{\small(1.67)}$ $41.76_{\small(1.12)}$ $39.62_{\small(1.32)}$ $33.88_{\small(2.03)}$ $51.65_{\small(1.55)}$ $41.78_{\small(1.27)}$ $42.14_{\small(2.71)}$
Wide&Deep $28.52_{\small(1.61)}$ $54.08_{\small(1.97)}$ $21.61_{\small(0.08)}$ $37.59_{\small(2.31)}$ $55.28_{\small(2.73)}$ $25.17_{\small(1.96)}$ $47.46_{\small(6.78)}$ $56.66_{\small(4.16)}$ $32.58_{\small(5.04)}$
DeepFM $29.56_{\small(0.13)}$ $51.16_{\small(0.65)}$ $18.70_{\small(0.79)}$ $35.11_{\small(1.25)}$ $51.43_{\small(0.57)}$ $21.20_{\small(0.70)}$ $43.21_{\small(2.78)}$ $52.15_{\small(0.18)}$ $27.87_{\small(2.33)}$
AFN $22.05_{\small(7.14)}$ $49.70_{\small(3.54)}$ $14.62_{\small(6.44)}$ $31.95_{\small(4.14)}$ $50.63_{\small(2.64)}$ $18.56_{\small(4.52)}$ $42.10_{\small(0.66)}$ $52.38_{\small(0.80)}$ $25.73_{\small(2.77)}$
MAMO $40.16_{\small(1.15)}$ $27.52_{\small(1.94)}$ $30.62_{\small(1.23)}$ $54.44_{\small(0.70)}$ $32.44_{\small(1.84)}$ $44.19_{\small(0.91)}$ $71.19_{\small(0.46)}$ $38.95_{\small(1.50)}$ $64.17_{\small(0.65)}$
TaNP $41.18_{\small(0.30)}$ $85.04_{\small(1.45)}$ $33.38_{\small(0.39)}$ $45.98_{\small(0.54)}$ $86.01_{\small(1.10)}$ $36.26_{\small(1.31)}$ $52.88_{\small(2.10)}$ $87.12_{\small(0.72)}$ $41.60_{\small(3.29)}$
MeLU $46.51_{\small(4.16)}$ $58.60_{\small(7.09)}$ $35.34_{\small(4.40)}$ $52.82_{\small(7.79)}$ $60.79_{\small(7.37)}$ $40.79_{\small(6.62)}$ $67.01_{\small(4.42)}$ $64.66_{\small(7.98)}$ $55.58_{\small(4.92)}$
HIRE $\mathbf{57.13_{\small(5.30)}}$ $\mathbf{89.31_{\small(2.91)}}$ $\mathbf{50.79_{\small(7.51)}}$ $\mathbf{64.05_{\small(2.57)}}$ $\mathbf{90.58_{\small(2.67)}}$ $\mathbf{56.40_{\small(5.44)}}$ $\mathbf{71.91_{\small(2.84)}}$ $\mathbf{92.95_{\small(2.34)}}$ $\mathbf{64.75_{\small(1.51)}}$
IC Popularity $42.19_{\small(6.50)}$ $84.29_{\small(0.84)}$ $30.90_{\small(5.72)}$ $44.37_{\small(4.29)}$ $85.05_{\small(0.35)}$ $32.51_{\small(4.52)}$ $48.04_{\small(3.13)}$ $85.69_{\small(0.06)}$ $33.29_{\small(3.94)}$
TransGNN $27.50_{\small(3.53)}$ $27.67_{\small(3.78)}$ $26.19_{\small(1.76)}$ $28.03_{\small(0.75)}$ $28.39_{\small(1.33)}$ $26.25_{\small(0.84)}$ $29.28_{\small(1.01)}$ $29.30_{\small(1.07)}$ $26.50_{\small(0.58)}$
NeuMF $40.70_{\small(0.66)}$ $36.32_{\small(4.07)}$ $32.82_{\small(4.14)}$ $51.42_{\small(2.25)}$ $38.13_{\small(4.22)}$ $43.03_{\small(7.73)}$ $59.30_{\small(8.68)}$ $41.03_{\small(4.30)}$ $57.29_{\small(2.00)}$
Wide&Deep $50.07_{\small(7.86)}$ $80.14_{\small(0.95)}$ $38.14_{\small(9.35)}$ $54.51_{\small(8.85)}$ $80.59_{\small(0.96)}$ $39.53_{\small(8.74)}$ $58.70_{\small(9.14)}$ $80.76_{\small(0.39)}$ $43.93_{\small(9.28)}$
DeepFM $52.46_{\small(8.13)}$ $81.10_{\small(1.62)}$ $40.92_{\small(7.72)}$ $55.50_{\small(2.12)}$ $81.56_{\small(1.81)}$ $41.21_{\small(9.28)}$ $59.50_{\small(2.13)}$ $82.00_{\small(2.45)}$ $44.37_{\small(9.95)}$
AFN $49.15_{\small(7.10)}$ $80.18_{\small(1.78)}$ $40.40_{\small(9.08)}$ $53.95_{\small(9.98)}$ $80.83_{\small(1.87)}$ $42.00_{\small(1.41)}$ $56.79_{\small(6.60)}$ $81.29_{\small(1.87)}$ $43.99_{\small(8.91)}$
MAMO $41.29_{\small(0.62)}$ $28.10_{\small(0.49)}$ $32.46_{\small(0.51)}$ $54.62_{\small(0.41)}$ $32.40_{\small(0.36)}$ $44.98_{\small(0.71)}$ $71.60_{\small(0.36)}$ $39.42_{\small(0.86)}$ $64.18_{\small(0.19)}$
TaNP $41.16_{\small(3.15)}$ $85.45_{\small(1.32)}$ $31.25_{\small(1.52)}$ $45.06_{\small(1.74)}$ $86.18_{\small(1.29)}$ $32.87_{\small(2.90)}$ $51.94_{\small(4.35)}$ $87.13_{\small(1.22)}$ $38.48_{\small(5.04)}$
MeLU $49.25_{\small(5.77)}$ $61.59_{\small(2.89)}$ $37.64_{\small(5.17)}$ $52.04_{\small(8.20)}$ $62.58_{\small(2.45)}$ $39.15_{\small(6.39)}$ $64.15_{\small(8.72)}$ $65.65_{\small(3.62)}$ $53.41_{\small(7.89)}$
HIRE $\mathbf{58.37_{\small(3.38)}}$ $\mathbf{89.25_{\small(0.56)}}$ $\mathbf{51.74_{\small(3.27)}}$ $\mathbf{64.18_{\small(2.44)}}$ $\mathbf{90.57_{\small(0.57)}}$ $\mathbf{56.49_{\small(2.06)}}$ $\mathbf{72.20_{\small(4.71)}}$ $\mathbf{92.62_{\small(0.65)}}$ $\mathbf{65.19_{\small(4.97)}}$
U&I C Popularity $26.47_{\small(1.39)}$ $84.46_{\small(2.53)}$ $16.55_{\small(1.57)}$ $29.57_{\small(1.58)}$ $84.63_{\small(1.89)}$ $18.02_{\small(1.60)}$ $34.48_{\small(1.03)}$ $85.13_{\small(1.19)}$ $21.51_{\small(1.35)}$
TransGNN $34.00_{\small(2.82)}$ $33.07_{\small(4.58)}$ $26.50_{\small(2.68)}$ $36.57_{\small(0.81)}$ $35.30_{\small(2.42)}$ $27.02_{\small(1.72)}$ $38.01_{\small(5.65)}$ $36.24_{\small(1.73)}$ $27.26_{\small(1.05)}$
NeuMF $38.29_{\small(0.97)}$ $42.21_{\small(3.29)}$ $29.76_{\small(1.05)}$ $43.42_{\small(1.20)}$ $43.04_{\small(3.45)}$ $32.74_{\small(2.29)}$ $49.22_{\small(0.74)}$ $43.64_{\small(2.05)}$ $36.61_{\small(1.70)}$
Wide&Deep $40.37_{\small(8.41)}$ $69.67_{\small(3.77)}$ $33.04_{\small(6.96)}$ $50.23_{\small(6.82)}$ $71.72_{\small(8.40)}$ $36.59_{\small(6.91)}$ $56.49_{\small(6.54)}$ $72.04_{\small(7.90)}$ $41.26_{\small(8.67)}$
DeepFM $39.27_{\small(7.71)}$ $68.48_{\small(7.31)}$ $30.18_{\small(7.10)}$ $47.50_{\small(2.94)}$ $69.00_{\small(8.38)}$ $36.49_{\small(4.29)}$ $54.14_{\small(9.65)}$ $72.66_{\small(5.03)}$ $38.81_{\small(9.81)}$
AFN $34.76_{\small(3.96)}$ $60.94_{\small(2.88)}$ $26.90_{\small(2.93)}$ $35.05_{\small(1.93)}$ $62.98_{\small(4.68)}$ $27.00_{\small(0.09)}$ $39.88_{\small(3.17)}$ $64.58_{\small(4.79)}$ $28.39_{\small(2.89)}$
MAMO $41.00_{\small(0.87)}$ $32.56_{\small(6.32)}$ $30.26_{\small(3.76)}$ $51.67_{\small(7.63)}$ $36.99_{\small(5.34)}$ $41.01_{\small(9.34)}$ $71.78_{\small(1.54)}$ $44.85_{\small(7.24)}$ $63.80_{\small(3.91)}$
TaNP $51.14_{\small(1.74)}$ $86.20_{\small(7.90)}$ $43.65_{\small(0.83)}$ $51.53_{\small(2.14)}$ $87.74_{\small(4.47)}$ $44.62_{\small(4.77)}$ $54.84_{\small(6.76)}$ $87.80_{\small(2.50)}$ $45.31_{\small(8.56)}$
MeLU $43.35_{\small(4.36)}$ $54.65_{\small(1.46)}$ $33.49_{\small(2.93)}$ $49.92_{\small(9.28)}$ $57.16_{\small(0.80)}$ $37.84_{\small(7.88)}$ $63.73_{\small(9.18)}$ $59.93_{\small(3.51)}$ $53.50_{\small(9.99)}$
HIRE $\mathbf{60.77_{\small(3.33)}}$ $\mathbf{90.60_{\small(0.25)}}$ $\mathbf{55.29_{\small(3.22)}}$ $\mathbf{69.03_{\small(4.30)}}$ $\mathbf{92.09_{\small(5.10)}}$ $\mathbf{62.84_{\small(4.05)}}$ $\mathbf{76.36_{\small(6.89)}}$ $\mathbf{93.92_{\small(0.64)}}$ $\mathbf{70.15_{\small(8.12)}}$
VI Experimental Studies
In this section, we give the test setting (§ VI-A) and report our substantial experimental results in the following facets: ① Compare the effectiveness of HIRE with the state-of-the-art approaches under 3 cold-start scenarios (§ VI-B). ② Compare the test efficiency of HIRE with the baseline approaches (§ VI-C). ③ Investigate the sensitivity of HIRE regarding key hyper-parameter configurations (§ VI-D). ④ Study the influence of different attention layers via an ablation study (§ VI-E). ⑤ Conduct a case study for HIRE on a movie rating task. (§ VI-F).
VI-A Experimental Setup
Datasets: We use 3 widely used datasets to evaluate HIRE and the baseline approaches, whose profiles are summarized in Table I. $\mathsf{MovieLens}$-$\mathsf{1M}$ [55] contains the ratings of users for movies, and the users and movies are associated with rich attributes. Following the previous following the previous literature [22], We randomly split 80% of users for training and validation, and 20% for test. movies released before 1997 are used for training and validation while the remaining are used for test, resulting in an approximate ratio of 8:2. $\mathsf{Bookcrossing}$ [56] is collected from Book-Crossing, containing ratings of users on books. $\mathsf{Douban}$ [57, 58] contains the ratings of users on musics, which are extracted from a rating website Douban. This dataset also contains friendship relations between users. Since users/items do not have attributes, we use the embeddings of the user/item ID as the attributes, which are generated by learnable linear weights. For $\mathsf{Bookcrossing}$ and $\mathsf{Douban}$, we randomly split $70\%$ of users/items into the train and validation set and $30\%$ into the test set, respectively. For all the datasets, train and validate sets are further split randomly by a ratio of 7:1. The test users/items are regarded as cold users/items where 10% of the interactions, i.e., no more than 3 interactions, in the datasets can be exploited by the recommendation models and the remaining 90% are used for model evaluation.
Baselines: To comprehensively evaluate HIRE, we compare with the following 12 baselines which fell into 6 categories:Basic Strategy: Popularity predicts ratings by summing the number of interactions and normalizing for each item. ❷ CF-based baselines: NeuMF [8], Wide&Deep [24], DeepFM [25] and AFN [26] adopt different neural network layers to model the interactions from user-item pairs. ❸ Social recommendation: GraphRec [15] models the users and items respectively by aggregating their local interactions via Graph Neural Network. Aggregation of the social relation among users is used to enhance user embeddings. We only compare GraphRec on $\mathsf{Douban}$ where the social relationship between users is available. ❹ HIN-based baselines: GraphHINGE [33] aggregates rich interactive patterns from an HIN. MetaHIN [36] incorporates multifaceted semantic contexts induced by meta-paths. We test MetaHIN and GraphHINGE only on $\mathsf{MovieLens}$-$\mathsf{1M}$ which provides sufficient attributes for constructing an HIN. Specifically, we define user, movie, genre, occupation, and age as different node types and create links between nodes, such as movie-genre interaction, user-age interaction, and user-movie interaction. For the other two datasets $\mathsf{Bookcrossing}$ and $\mathsf{Douban}$, due to lack of sufficient attributes of user/item, the two baselines cannot be applied to them. ❺ Meta-learning baselines: MAMO [21], MeLU [22] and TaNP [20] adopt meta-learning for cold-start recommendation in a multi-task fashion. MAMO and MeLU support all the 3 cold-start scenarios while TaNP only supports user cold-start originally. Here, we extend the task settings of TaNP to item cold-start and user & item cold-start. ❻ Transformer-based baselines: TransGNN [44] integrates transformer and GNN layers to mutually enhance their capabilities.
Implementation details: We introduce the model and training details of HIRE as well as the baselines. For HIRE, the model is equipped with $3$ HIM blocks in which each MHSA layer has $8$ heads and the hidden dimension of each head is set to $16$. The model is trained by a LAMB optimizer [59] with $\beta=(0.9,0.999)$ and $\epsilon=10^{-6}$, and a Lookahead [60] wrapper with slow update rate $\alpha=0.5$ and $k=6$ steps between updates. We use a flat-then-anneal learning rate scheduler which flats at the base learning rate for $70\%$ of steps, and then anneals following a cosine schedule to 00 by the end of training. We set the base learning rate to $10^{-3}$ and the gradient clip threshold to $1.0$. The optimizer and the scheduler are widely used for training language models composed of deep MHSA layers. Regarding the prediction contexts, we set the number of users and items in each context to 32 by default. In the training and test, $90\%$ of observed ratings are masked for prediction and the remaining $10\%$ are associated with the context input. The learning framework of HIRE is built on PyTorch.
For the baseline approaches, we use their released source code and keep the hyper-parameters as their default settings. All the models are trained until convergence. For the four CF-based approaches, GraphRec and GraphHINGE, to achieve a fair comparison, we use all the observed user-item ratings in the sampled training contexts, together with the $10\%$ unmasked user-item ratings in the test context as the ground truth to train the models. We use the prediction context to test TransGNN, where a few number of user-item interaction is used for training and the remaining is used for test. For meta-learning approaches, in each task, $10\%$ of the observed ratings are used as the ground truth in the support set and the remaining $90\%$ are used in the query set. Training and testing are conducted on one Tesla V100 with 16GB memory.
Evaluation Metrics: We adopt 3 commonly used metrics following the previous work [20] to evaluate the recommendation performance in our testing, including: Precision, Mean Average Precision (MAP) and Normalized Discounted Cumulative Gain [61] (NDCG). Top $k$ actual rating values sorted by predicted rating values are used to calculate the above metrics, where $k\in\{5,7,10\}$.